# Credal Prediction based on Relative Likelihood

### v1

@@ B.7 Computing Resources
To run the experiments presented in this work, we utilized the computing resources detailed in Table 3. The total estimated GPU usage amounts to approximately $750$ hours.
Table 3: Specifications of Computing Resources
@@ C.1 Out-of-Distribution detection
In addition to the OoD experiments in Section 5, we evaluate a broad range of values $\alpha$ for the CreEnsα approach.
@@
$\text{CreEns}_{0.2}$ $0.950\scriptstyle{\pm 0.001}$ $0.909\scriptstyle{\pm 0.001}$ $0.911\scriptstyle{\pm 0.000}$ $0.946\scriptstyle{\pm 0.001}$ $0.885\scriptstyle{\pm 0.001}$
@@
CreBNN $0.907\scriptstyle{\pm 0.006}$ $0.885\scriptstyle{\pm 0.002}$ $0.880\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.002}$ $0.859\scriptstyle{\pm 0.002}$
@@ C.2.1 Tobias Initialization
We study the impact of the ToBias initialization constant $\beta$ on the coverage and efficiency of the resulting credal predictor on the ChaosNLI dataset. To do so, we vary the constant, taking values $\beta\in\{5,10,20,30,50,80,100,200,500\}$. The ensemble is then trained as proposed in Algorithm 1. The coverage and efficiency Pareto front is shown in Figure 5 with the mean and standard deviation over three runs. We split the results into three Figures for better readability.
@@ C.2.2 Number of Ensemble Members
We study the impact of the numbers of ensemble members $M$ on the coverage and efficiency of the resulting credal predictor on the ChaosNLI dataset. To do so, we vary the constant, taking values $M\in\{1,2,3,4,5,10,20,30,50\}$. The ensemble is then trained as proposed in Algorithm 1. The coverage and efficiency Pareto front is shown in Figure 6 with the mean and standard deviation over three runs. We split the results into three Figures for better readability.
@@
Table 5: Average entropy of ground-truth distribution in the test sets of ChaosNLI, CIFAR-10 and QualityMRI.
@@ C.3 Data Uncertainty
The datasets used in our experiments, namely ChaosNLI, CIFAR-10, and QualityMRI, differ in their inherent levels of (aleatoric) uncertainty. This is reflected, for example, in the degree of disagreement among the collected annotations, which serve as the ground-truth distributions in our evaluations. To provide further insight, we report the average entropy of these ground-truth distributions in Table 5.
@@ C.4 Performance of Ensemble Members
We visualized the trade-off between coverage and efficiency in Figure 4, and provide the corresponding numerical values in Table 6. We also report the accuracies of the individual predictors within each ensemble for both the baselines and our proposed method to provide insight into their standalone performance (cf. Tables 7 and 8). For CreNet, whose ensemble members directly predict probability intervals, we use their intersection probability technique to derive pointwise predictions [Wang et al., 2024a].
@@
CreNet $0.001\scriptstyle{\pm 0.002}$ $0.978\scriptstyle{\pm 0.008}$ $0.094\scriptstyle{\pm 0.010}$ $0.999\scriptstyle{\pm 0.000}$ $0.188\scriptstyle{\pm 0.020}$ $0.792\scriptstyle{\pm 0.003}$
CreBNN $0.195\scriptstyle{\pm 0.019}$ $0.648\scriptstyle{\pm 0.019}$ $0.331\scriptstyle{\pm 0.004}$ $0.871\scriptstyle{\pm 0.003}$ $0.780\scriptstyle{\pm 0.312}$ $0.184\scriptstyle{\pm 0.255}$
CreEns0.0 $0.298\scriptstyle{\pm 0.041}$ $0.607\scriptstyle{\pm 0.037}$ $0.000\scriptstyle{\pm 0.000}$ $0.963\scriptstyle{\pm 0.000}$ $0.329\scriptstyle{\pm 0.040}$ $0.611\scriptstyle{\pm 0.023}$
CreEns0.2 $0.165\scriptstyle{\pm 0.007}$ $0.769\scriptstyle{\pm 0.010}$ $0.000\scriptstyle{\pm 0.000}$ $0.983\scriptstyle{\pm 0.000}$ $0.177\scriptstyle{\pm 0.060}$ $0.787\scriptstyle{\pm 0.020}$
CreEns0.4 $0.096\scriptstyle{\pm 0.008}$ $0.834\scriptstyle{\pm 0.006}$ $0.000\scriptstyle{\pm 0.000}$ $0.986\scriptstyle{\pm 0.000}$ $0.118\scriptstyle{\pm 0.059}$ $0.844\scriptstyle{\pm 0.021}$
CreEns0.6 $0.043\scriptstyle{\pm 0.003}$ $0.886\scriptstyle{\pm 0.005}$ $0.000\scriptstyle{\pm 0.000}$ $0.988\scriptstyle{\pm 0.000}$ $0.081\scriptstyle{\pm 0.057}$ $0.899\scriptstyle{\pm 0.012}$
CreEns0.8 $0.012\scriptstyle{\pm 0.002}$ $0.941\scriptstyle{\pm 0.002}$ $0.000\scriptstyle{\pm 0.000}$ $0.993\scriptstyle{\pm 0.000}$ $0.054\scriptstyle{\pm 0.033}$ $0.950\scriptstyle{\pm 0.008}$
CreEns0.9 $0.000\scriptstyle{\pm 0.000}$ $0.976\scriptstyle{\pm 0.001}$ $0.000\scriptstyle{\pm 0.000}$ $0.997\scriptstyle{\pm 0.000}$ $0.016\scriptstyle{\pm 0.023}$ $0.983\scriptstyle{\pm 0.002}$

### v2

@@ B.7 Computing Resources
To run the experiments presented in this work, we utilized the computing resources detailed in Table˜3. The total estimated GPU usage amounts to approximately $750$ hours.
Table 3: Specifications of Computing Resources.
@@ C.1 Out-of-Distribution detection
In addition to the OoD experiments in Section˜5, we evaluate a broad range of values $\alpha$ for the CreEns approach.
@@
$\text{CreEns}_{0.2}$ $50\scriptstyle{\pm 0.001} 0.909$$09\scriptstyle{\pm 0.001}$ $0.911\scriptstyle{\pm 0.000}$ $0.946\scriptstyle{\pm 0.001}$ $0.885\scriptstyle{\pm 0.001}$
@@
CreBNN $0.907\scriptstyle{\pm 0.006}$ $85\scriptstyle{\pm 0.002} 0.880$$80\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.002}$ $0.859\scriptstyle{\pm 0.002}$
@@ C.2.1 ToBias Initialization
We study the impact of the ToBias initialization constant $\beta$ on the coverage and efficiency of the resulting credal predictor on the ChaosNLI dataset. To do so, we vary the constant, taking values $\beta\in\{5,10,20,30,50,80,100,200,500\}$. The ensemble is then trained as proposed in Algorithm˜1. The coverage and efficiency Pareto front is shown in Figure˜5 with the mean and standard deviation over three runs. We split the results into three Figures for better readability.
@@ C.2.2 Number of Ensemble Members
We study the impact of the numbers of ensemble members $M$ on the coverage and efficiency of the resulting credal predictor on the ChaosNLI dataset. To do so, we vary the constant, taking values $M\in\{1,2,3,4,5,10,20,30,50\}$. The ensemble is then trained as proposed in Algorithm˜1. The coverage and efficiency Pareto front is shown in Figure˜6 with the mean and standard deviation over three runs. We split the results into three Figures for better readability.
@@
Table 5: Average entropy of ground-truth distributions in the test sets of ChaosNLI, CIFAR-10 and QualityMRI.
@@ C.3 Data Uncertainty
The datasets used in our experiments, namely ChaosNLI, CIFAR-10, and QualityMRI, differ in their inherent levels of (aleatoric) uncertainty. This is reflected, for example, in the degree of disagreement among the collected annotations, which serve as the ground-truth distributions in our evaluations. To provide further insight, we report the average entropy of these ground-truth distributions in Table˜5.
@@ C.4 Performance of Ensemble Members
We visualized the trade-off between coverage and efficiency in Figure˜4, and provide the corresponding numerical values in Table˜6. We also report the accuracies of the individual predictors within each ensemble for both the baselines and our proposed method to provide insight into their standalone performance (cf. Tables˜7 and 8). For CreNet, whose ensemble members directly predict probability intervals, we use their intersection probability technique to derive pointwise predictions [wangCredalDeep2024].
@@
CreNet $0.001\scriptstyle{\pm 0.002}$ $978\scriptstyle{\pm 0.008} 0.0$$094\scriptstyle{\pm 0.010}$ $0.999\scriptstyle{\pm 0.000}$ $0.188\scriptstyle{\pm 0.020}$ $0.792\scriptstyle{\pm 0.003}$
CreBNN $0.195\scriptstyle{\pm 0.018}$ $0.649\scriptstyle{\pm 0.017}$ $0.331\scriptstyle{\pm 0.005}$ $0.871\scriptstyle{\pm 0.003}$ $898\scriptstyle{\pm 0.133} 0.0$$090\scriptstyle{\pm 0.124}$
CreEns0.0 $0.298\scriptstyle{\pm 0.041}$ $607\scriptstyle{\pm 0.037} 0.0$$000\scriptstyle{\pm 0.000}$ $0.963\scriptstyle{\pm 0.000}$ $0.329\scriptstyle{\pm 0.040}$ $0.611\scriptstyle{\pm 0.023}$
CreEns0.2 $0.165\scriptstyle{\pm 0.007}$ $769\scriptstyle{\pm 0.010} 0.0$$000\scriptstyle{\pm 0.000}$ $0.983\scriptstyle{\pm 0.000}$ $0.177\scriptstyle{\pm 0.060}$ $0.787\scriptstyle{\pm 0.020}$
CreEns0.4 $0.096\scriptstyle{\pm 0.008}$ $834\scriptstyle{\pm 0.006} 0.0$$000\scriptstyle{\pm 0.000}$ $0.986\scriptstyle{\pm 0.000}$ $0.118\scriptstyle{\pm 0.059}$ $0.844\scriptstyle{\pm 0.021}$
CreEns0.6 $0.043\scriptstyle{\pm 0.003}$ $886\scriptstyle{\pm 0.005} 0.0$$000\scriptstyle{\pm 0.000}$ $988\scriptstyle{\pm 0.000} 0.0$$081\scriptstyle{\pm 0.057}$ $0.899\scriptstyle{\pm 0.012}$
CreEns0.8 $0.012\scriptstyle{\pm 0.002}$ $941\scriptstyle{\pm 0.002} 0.0$$000\scriptstyle{\pm 0.000}$ $993\scriptstyle{\pm 0.000} 0.0$$054\scriptstyle{\pm 0.033}$ $0.950\scriptstyle{\pm 0.008}$
CreEns0.9 $0.000\scriptstyle{\pm 0.000}$ $976\scriptstyle{\pm 0.001} 0.0$$000\scriptstyle{\pm 0.000}$ $997\scriptstyle{\pm 0.000} 0.0$$016\scriptstyle{\pm 0.023}$ $0.983\scriptstyle{\pm 0.002}$